7) a) If the corporate tax rate is 40% and the individual tax rate is 20% (and these are the only taxes), if a corporation has $50,000 of operating income, what is the after tax income for its owners? b) The tax rate is 15% between $0-$50,000 and 25% between $50,001-$100,000. If you earn $60,000, then what is your average and marginal tax rates? A bank has quoted to you a monthly rate of 0.25% on a short-term loan. what is the stated annual rate of that loan? c) d) What is the future value of S8 in three years if the interest rate is 20% and the compounding is continuous? e What is the future value of a 3-year annuity due with $50 cash flows if the annually compounded interest rate is 6%? f) If we need to find the future value of a current cash flow and interest is compounded quarterly, how many periods of compounding occur in two and half years? What is the difference between the present value of a 3-year ordinary annuity with a $20 cash flow and the present value of a 3-year annuity due with a $20 cash flow when the interest rate is 3.5%?

Need assitance with E, F, G. Please refrain from using Excel! Thanks!
Answer #1
7)
e) Future Value of annuity = Annual cash flow * FVIFA(interest rate, n)
Future Value of annuity = 50 * FVIFA(6%, 3) = 50 * 3.1836 = 159.18
f) The compounding periods would be 2.5 years / 0.25 year = 10 periods
g) Present value of 3 year ordinary annuity has no discounting factor whereas
Present value of 3 year annuity when the interest rate is 3.5% has a discount
factor of 3.5%. The present value in ordinary annuity is the total of all annuities whereas
the annuity with interest rate got discounted with discount factors and final
value got reduced of sum total of all annuities. Present value of ordinary annuity = 20*3=60
whereas present value with interest rate = 20 * PVIFA(3.5%,3) = 20 *2.8016 = 56.03